# Random Latent Exploration for Deep Reinforcement Learning

### v1

@@ Abstract
The ability to efficiently explore high-dimensional state spaces is essential for the practical success of deep Reinforcement Learning (RL). This paper introduces a new exploration technique called Random Latent Exploration (RLE), that combines the strengths of bonus-based and noise-based (two popular approaches for effective exploration in deep RL) exploration strategies. RLE leverages the idea of perturbing rewards by adding structured random rewards to the original task rewards in certain (random) states of the environment, to encourage the agent to explore the environment during training. RLE is straightforward to implement and performs well in practice. To demonstrate the practical effectiveness of RLE, we evaluate it on the challenging Atari and IsaacGym benchmarks and show that RLE exhibits higher overall scores across all the tasks than other approaches.
@@ 1 Introduction
Instead, we hypothesize that it is easier to explore by training the agent to achieve a variety of diverse goals, i.e. exploring the space of goals. Following this intuition, in this work, we introduce a different exploration strategy where instead of injecting noise or adding bonuses to encourage exploration, we train the agent to achieve various goals from a goal space $\mathcal{Z}$. Our key intuition is that if the goals are designed such that different goals incentivize the agent to explore diverse parts of the environment, then training an agent to achieve these goals will lead it to explore a variety of states, eventually leading it to find and exploit high task rewards. Unfortunately, representing and selecting "the right goals" is challenging in practice, as they may be dependent on the underlying task in complex ways that the learner cannot anticipate in advance.
To circumvent the issue of finding the right goals to target, we propose Random Latent Exploration (RLE), where the agent’s policy is conditioned on random vectors sampled from a fixed distribution as goals, which define additional reward bonuses differing at each state based on the chosen vectors. In particular, the random vectors act as latent goals, each creating different reward functions that encourage the agent to reach different parts of the environment. By sampling enough random vectors during training, the agent is trained to pursue many different goals, thus resulting in deep exploration. Our experiments show that RLE leads to significantly more diverse and deeper trajectories than traditional noise-based methods, without the need for complex bonus computation. This makes RLE both easy to implement and effective in practice.
To show the effectiveness of our approach, we evaluated RLE in Atari—a popular discrete action space deep RL benchmark (Bellemare et al., 2013), and IsaacGym—a popular continuous control deep RL benchmark (Makoviychuk et al., 2021), each consisting of many different tasks with varying degrees of exploration difficulty. We implement our method on top of the popular RL algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017) and compare it with PPO in other exploration strategies. Our experimental results demonstrate that RLE improves over standard PPO in Atari and IsaacGym. Furthermore, RLE also exhibits a higher aggregated score across all tasks in Atari than other exploration methods, including RND (Burda et al., 2019) and randomized value function strategies (Fortunato et al., 2017). Importantly, these improvements were obtained by simply adding RLE on the top of the base PPO implementation while only changing the learning rate (and discount rate for Atari). We use the same hyperparameters across all 57 Atari games and the same hyperparameters across all 9 IsaacGym tasks, thus highlighting the generality of our approach as a plug-in utility.
@@ 2 Preliminaries
RL is a popular paradigm for solving sequential decision-making problems (Sutton & Barto, 2018) where an agent operates in an unknown environment (Sutton & Barto, 2018) and aims to improve its performance through repeated interactions with the environment. At each round of interaction, the agent starts from an initial state $s_{0}$ of the environment and collects a trajectory. At each timestep $t$ within that trajectory, the agent perceives the state $s_{t}$, takes an action $a_{t}\sim\pi(.|s_{t})$ with its policy $\pi$, receives a task reward $r_{t}=r(s_{t},a_{t})$, and transitions to a next state $s_{t+1}$ until reaching terminal states, after which a new trajectory is initialized from $s_{0}$ and the above repeats. The goal of the agent is to learn a policy $\pi$ that maximizes expected return $\mathbb{E}_{\pi}\Big{[}\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})\Big{]}$ in a trajectory. A straightforward approach is to estimate the expected return of a policy by rolling out trajectories $(s_{0},a_{0},s_{1},\cdots,)$ through Monte Carlo sampling (Konda & Tsitsiklis, 1999), and then optimizing this to find the optimal policy, but unfortunately, the corresponding estimates are of high variance and thus often require a huge number of data. Thus, in practice, various RL algorithms learn a value function (or value network) $V^{\pi}$ from the interaction that approximates
@@
Exploration. As the reward may be delayed and not immediately presented to the agent, the agent may need to take many actions and visit a sequence of states without rewards before it receives any learning signal (reward). As such, taking greedy action $a_{t}$ at each step that maximizes immediate reward $r(s_{t},a_{t})$ does not necessarily lead to a high return. Thus, RL algorithms require “exploring” states and actions that may lead to low immediate rewards but could potentially end up with high return in the long run. We refer to this process as exploration throughout this paper.
@@ 3 Our Method: Random Latent Exploration
Problem statement. Our goal in this paper is to create an exploration strategy that improves the standard action noise exploration method overall on the majority of tasks and domains (both discrete and continuous control) and yet is easy to implement. Toward that end, we provide a simple exploration method that combines the best of both worlds and bridges the approach of using reward bonuses (Bellemare et al., 2016; Burda et al., 2019) with noises or posterior sampling for exploration (Osband et al., 2016a, b; Fortunato et al., 2017).
Challenge: Ensuring diversity. The key challenge in noise-based exploration methods lies in their ability to generate diverse behaviors within the environment. Specifically, if noise-perturbed policies bear too much similarity, they will yield nearly identical trajectories $(s_{0},a_{0},r_{0},s_{1},\cdots)$, consequently limiting exploration to a narrow region of the environment. The primary technical challenge, then, is devising a method that ensures such diversity in the generated trajectories.
Approach: Randomized rewards. To overcome the above challenge, we propose enhancing trajectory diversity by altering the agent’s rewards, inspired by skill discovery methods (Eysenbach et al., 2018) that use varied reward functions to develop diverse skills resulting in different state visitation distribution. In particular, in every episode of training, we randomly perturb the given task reward function by adding a randomized reward to certain (random) parts of the state space. Then, we train policies to maximize the composite of randomized and task rewards. The key idea is that in every round of interaction, the random rewards added to certain random parts of the state space will incentivize the agent to visit those areas, and if these random areas are diverse enough, we will get diverse behaviors in the environment during training—thus incentivizing exploration. However, since the random rewards are repeatedly resampled and thus keep on changing during training to ensure stable and effective learning, both policies and value functions must be aware of the specific random reward function in use; otherwise, the changing reward functions will comprise a partially observable MDP (Kaelbling et al., 1998). We take inspiration from the Universal Value Function Approximator (UVFA) (Schaul et al., 2015), which trains networks based on varying goals. We adopt their approach by equating their goals to different reward functions, thus making the policy $\pi$ and the value function $V^{\pi}$ condition on the sampled reward function. This ensures that the random rewards no longer appear as noises to the policy. The remaining questions are:
@@
We outline our implementation of the above idea in the next section and defer full implementation details to Appendix B.
@@ 3.1 Algorithmic Implementation
Randomized reward functions. An effective approach to implementing randomized reward functions necessitates adherence to two fundamental principles. First, the randomized rewards require correlating with states; otherwise, they would appear as white noises, which does not help exploration as shown by Fortunato et al. (2017). Second, the function must be factorized to allow both the policy and the value function to condition the factors of the randomized reward function. This avoids the perception of the reward as partially observable by the policy. To fulfill these criteria, we implement the randomized reward function (denoted as $F$) as the dot product of the state feature and a randomly chosen latent vector:
@@
where $\phi:\mathcal{S}\to\mathbb{R}^{d}$ is a feature extractor that transforms a state into a $d$-dimensional vector, and $\boldsymbol{z}\in\mathbb{R}^{d}$ represents a latent vector. Randomized rewards for each state are generated by sampling $\boldsymbol{z}$ from a given distribution $P_{\boldsymbol{z}}$, and then setting rewards as $F(s,\boldsymbol{z})$.
Latent conditioned policy and value network. Recall that the policy and the value function have to be aware of the state and the random variable that factorizes the randomized reward function $F$. To achieve this, we augment the input to the policy and the value functions with the latent vector $\boldsymbol{z}$. The resulting policy and the value networks are $\pi(.|s,\boldsymbol{z})$ and $V^{\pi}(s,\boldsymbol{z})$. We train the latent-conditioned value network to approximate the expected sum of the original reward and the randomized rewards as below
$V^{\pi}(s,\boldsymbol{z})\approx\mathbb{E}_{\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t})+F(s_{t+1},\boldsymbol{z}))\right],$ (3)
and train the latent-conditioned policy $\pi$ to maximize $V^{\pi}(s,\boldsymbol{z})$ at every state $s$ and latent vector $\boldsymbol{z}$. Both value and policy networks can be trained with any off-the-shelf RL algorithms, e.g. PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), A3C (Mnih et al., 2016), SAC (Haarnoja et al., 2018).
Latent vector sampling. To randomize the latent vector $\boldsymbol{z}$, rather than resampling it at each timestep, which could cause dithering actions due to fast-changing latent vectors fed to the policy, it is resampled at the start of each trajectory. This ensures each trajectory is rolled out under the same policy and latent vector $\boldsymbol{z}$, maintaining temporal consistency crucial for deep exploration, as indicated by prior studies (Osband et al., 2016a; Fortunato et al., 2017). The distribution of $\boldsymbol{z}$ will be elaborated in Section 4.
As we train the policy conditioned on the latent factor of the randomly sampled reward function, we term our method as Random Latent Exploration (RLE). We outline the algorithm in Algorithm 1 and present the detailed version in Algorithm 2. Note that at line 6 in Algorithm 1, we compute the randomized reward at the next state $s_{t+1}$ since the next state reflects the credit of taking action $a_{t}$ and state $s_{t}$. This choice is also common in prior works that compute exploration bonuses at each step (Burda et al., 2019).
@@ 4 Experiments
We aim to show that our RLE improves over the action noise exploration method typically used in many RL algorithms (Schulman et al., 2017; Mnih et al., 2015) on the majority of tasks in both discrete and continuous control domains. Throughout the experiments, we train the agent for each task using PPO (Schulman et al., 2017) since it is a popular RL algorithm used in both discrete and continuous control tasks. Standard PPO implementation (Schulman et al., 2017) explores by sampling actions from the policy (i.e., a Boltzmann distribution over actions) learned.
@@
  • NoisyNet (Fortunato et al., 2017): We chose it to be the representative baseline from the family of noise-based exploration (Osband et al., 2016a; Fortunato et al., 2017; Plappert et al., 2017) because it has been used in prior works on benchmarking exploration strategies (Chen et al., 2022; Taïga et al., 2019).
@@ 4.1 Illustrative Experiments on FourRoom
To test whether our method can perform deep exploration that is essential in many tasks requiring exploration, we start by running toy experiments on the the FourRoom environment (Sutton et al., 1999). We explain the setting of FourRoom environment and the results as follows.
Setup. Figure 1 illustrates FourRoom environment with $50\times 50$ states. The environment consists of four rooms separated by solid walls (which the agent can’t cross) and connected with small openings of a single state each (which the agent needs to go through to travel across rooms).
Figure 1: FourRoom environment. The agent starts at the top-right state (denoted by red ’S’) and can move left, right, up, and down. The black bars denote walls that block the agent’s movement.
@@
The agent perceives the $(x,y)$ coordinates as the state input, and can take an action to move left, right, up, and down (if not interrupted by a wall). At the beginning of each trajectory, the agent always starts from the top-right corner of the room (denoted by the letter “S”). In this study, we always give zero rewards to the agent since we are interested in understanding and comparing how each exploration strategy behaves in the presence of sparse rewards, or even zero everywhere. This is also known as reward-free exploration.We also perform experiments on FourRoom with an sparse task reward of $1$ at the bottom-left corner. The results and visualization of visitation counts are deferred to Appendix C.
@@
We compared the agents trained with different exploration strategies: PPO, NoisyNet, RND, and RLE. The corresponding policy in each of these exploration strategies is trained with Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) for 2.5 million timesteps. For RLE, the feature extractor defined in Equation 2 is set to be a randomly initialized neural network with one hidden layer, the output layer of which has the same dimension as $\boldsymbol{z}\sim P_{\boldsymbol{z}}$. Further implementation details are available in Section B.2. We remark that, because of the walls, the FourRoom environment requires deep exploration to go to states distant from the initial states.
@@
How does the latent vector influence the generated trajectories? The exploration of RLE is driven by sampling the latent vector $\boldsymbol{z}$ to change the behavior of the policy network $\pi$ and the objective of the value network $V^{\pi}$. One may wonder what is the impact of $\boldsymbol{z}$ on the induced policy’s behaviors. To evaluate this, we sample different latent vector $\boldsymbol{z}\sim P_{z}$ and rollout trajectories with the policy conditioning on those $\boldsymbol{z}$, plotting each trajectory in a different color in Figure 2. For this plot, we chose the checkpoint of the policy network stored in the middle of training (i.e., 1.5 million timesteps) since we want to study the behaviors of the policy network before it converges. Figure 2 demonstrates that the trajectories generated by different latent vectors $\boldsymbol{z}$ can span across all four rooms. This shows that altering the latent vector $\boldsymbol{z}$ can produce diverse trajectories and that latent vectors can influence the behaviors of the policy network. We provide the rollouts of trajectories for random seeds for RLE, PPO, RND and NoisyNet in Appendix C.
@@
Explaining the observed trajectory diversity. To explain why the trajectories generated by an RLE policy are diverse, we plot several of the reward functions induced by sampling different $\boldsymbol{z}$ in Figure 10. These plots demonstrate the diversity of the random rewards, each of which can guide the policy to a different part of the state space.
State Visitations in the environment. We plot the state visitation counts of each exploration strategies and present the results in Figure 3. The results shows that PPO’s state visitation centers around the initial room (i.e, top-right), indicating that action noises are not able to bring the agent far away from the initial state. In contrast, we see that RLE, RND, and NoisyNet are all able to frequently reach the rooms beyond the initial room, with RLE visitation count spread out across the four rooms. This suggests that RLE is capable of doing deep exploration similar to prior deep exploration algorithms for this environment.

### v2

@@ Abstract
We introduce Random Latent Exploration (RLE), a simple yet effective exploration strategy in reinforcement learning (RL). On average, RLE outperforms noise-based methods, which perturb the agent’s actions, and bonus-based exploration, which rewards the agent for attempting novel behaviors. The core idea of RLE is to encourage the agent to explore different parts of the environment by pursuing randomly sampled goals in a latent space. RLE is as simple as noise-based methods, as it avoids complex bonus calculations but retains the deep exploration benefits of bonus-based methods. Our experiments show that RLE improves performance on average in both discrete (e.g., Atari) and continuous control tasks (e.g., Isaac Gym), enhancing exploration while remaining a simple and general plug-in for existing RL algorithms.
@@ 1 Introduction
Instead, we hypothesize that it is easier to explore by training the agent to achieve a variety of diverse goals, i.e., exploring the space of goals (Smith & Gasser, 2005; Forestier et al., 2022). We build upon this intuition and introduce an exploration strategy where instead of injecting noise or adding bonuses to encourage exploration, we train the agent to achieve various goals from a goal space $\mathcal{Z}$. Our key intuition is that if the goals are designed such that different goals incentivize the agent to explore diverse parts of the environment, then training an agent to achieve these goals will lead it to explore a variety of states, eventually leading it to find and achieve high task rewards. While goal-based exploration has been well studied (Nair et al., 2018; Ecoffet et al., 2019, 2021; Torne Villasevil et al., 2023), representing and selecting "the right goals" is challenging in practice, as they may be dependent on the underlying task in complex ways that the learner cannot anticipate in advance.
To circumvent the issue of finding the right goals to target, we propose Random Latent Exploration (RLE), where the agent’s policy is conditioned on random vectors sampled, $z\sim\mathcal{Z}$, from a fixed distribution as goals. Each goal ($z$) defines an exploration bonus at every state, and different goals define different bonuses. In particular, the random vectors act as latent goals, each creating different reward functions that encourage the agent to reach different parts of the environment. By sampling enough random vectors during training, the agent is trained to pursue many different goals, thus resulting in deep exploration. Our experiments show that RLE leads to significantly more diverse and deeper trajectories than traditional noise-based methods, without the need for complex bonus computation. This makes RLE both easy to implement and effective in practice.
To show the effectiveness of our approach, we evaluated RLE in Atari—a popular discrete action space deep RL benchmark (Bellemare et al., 2013), and IsaacGym—a popular continuous control deep RL benchmark (Makoviychuk et al., 2021), each consisting of many different tasks with varying degrees of exploration difficulty. We implement our method on top of the popular RL algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017) and compare it with PPO in other exploration strategies. Our experimental results demonstrate that RLE improves over standard PPO in Atari and IsaacGym. Furthermore, RLE also exhibits a higher aggregated score across all tasks in Atari than other exploration methods, including RND (Burda et al., 2019) and randomized value function strategies (Fortunato et al., 2017). Importantly, these improvements were obtained by simply adding RLE on the top of the base PPO implementation while only changing the learning rate (and discount rate for Atari). We use the same hyperparameters across all 57 Atari games and the same hyperparameters across all 9 IsaacGym tasks, highlighting the generality of our approach as a plug-in utility.
@@ 2 Preliminaries
RL is a popular paradigm for solving sequential decision-making problems (Sutton & Barto, 2018) where an agent operates in an unknown environment (Sutton & Barto, 2018) and aims to improve its performance through repeated interactions with the environment. At each round of interaction, the agent starts from an initial state $s_{0}$ of the environment and collects a trajectory. At each timestep $t$ within that trajectory, the agent perceives the state $s_{t}$, takes an action $a_{t}\sim\pi(.|s_{t})$ with its policy $\pi$, receives a task reward $r_{t}=r(s_{t},a_{t})$, and transitions to a next state $s_{t+1}$ until reaching terminal states, after which a new trajectory is initialized from $s_{0}$ and the above repeats. The goal of the agent is to learn a policy $\pi$ that maximizes expected return $\mathbb{E}_{\pi}\Big{[}\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})\Big{]}$. A straightforward approach is to estimate the expected return of a policy by rolling out trajectories $(s_{0},a_{0},s_{1},\cdots,)$ through Monte Carlo sampling (Konda & Tsitsiklis, 1999), and then optimizing this to find the optimal policy, but unfortunately, the corresponding estimates are of high variance and thus often require a huge number of data. Thus, in practice, various RL algorithms learn a value function (or value network) $V^{\pi}$ from the interaction that approximates
@@
Exploration. As the reward may be delayed and not immediately presented to the agent, the agent may need to take many actions and visit a sequence of states without rewards before it receives any learning signal (reward). As such, taking greedy action $a_{t}$ at each step that maximizes immediate reward $r(s_{t},a_{t})$ does not necessarily lead to a high return. Thus, RL algorithms require “exploring” states and actions that may lead to low immediate rewards but could potentially end up with high return in the long run. We refer to this process as exploration throughout this paper (Sutton & Barto, 2018).
@@ 3 Our Method: Random Latent Exploration
Problem statement. We aim to develop an easy-to-implement exploration strategy that improves over the standard action noise exploration in discrete and continuous action spaces.
Challenge: Ensuring diversity. The key challenge in noise-based exploration is their limited ability to generate diverse behaviors. Although noise-based exploration changes the agent’s actions at each timestep, in practice these perturbations are too local to deviate significantly from the usual trajectories. As a result, the generated trajectories can appear similar to each other, limiting exploration to a narrow region of the environment. The primary technical challenge, then, is devising a method that ensures such diversity in the generated trajectories.
Approach: Randomized rewards. To overcome the above challenge, we propose enhancing trajectory diversity by altering the agent’s rewards, inspired by skill discovery methods (Eysenbach et al., 2018) that learn skills based on different reward functions that push the agent to visit different parts of the state space for different skills. Following a similar idea, in every training episode, we perturb the given task reward function by adding a randomized state-dependent reward function, and we train policies to maximize the sum of randomized and task rewards. The key idea is that in every round of interaction, the agent is incentivized to visit areas of the state space with high randomized rewards, and if these random areas are diverse enough, we will get diverse behaviors in the environment during training—thus incentivizing exploration. However, since the random rewards are repeatedly resampled and thus keep on changing during training to ensure stable and effective learning, both policies and value functions must be aware of the specific random reward function in use; otherwise, the changing reward functions will comprise a partially observable MDP (Kaelbling et al., 1998). We take inspiration from the Universal Value Function Approximator (UVFA) (Schaul et al., 2015), which trains networks based on varying goals. We adopt their approach by equating their goals with different reward functions, thus making the policy $\pi$ and the value function $V^{\pi}$ condition on the sampled reward function. This ensures that the random rewards no longer appear to be noise to the policy. The remaining questions are:
@@
In the next section, we outline our implementation of the above idea, and the full implementation is provided in Appendix B.
@@ 3.1 Algorithmic Implementation
Randomized reward functions. We design a practical approach to implement randomized reward functions using two principles. First, the randomized reward function must depend on the observation; otherwise the rewards would be random noise, which does not help exploration, as shown by Fortunato et al. (2017). Second, the policy and value functions must be conditioned on the randomized reward function so that the reward function is fully observable to them. Following these criteria, we efficiently implement the randomized reward function as the dot product of the state feature and a randomly chosen latent vector:
@@
where $\phi:\mathcal{S}\to\mathbb{R}^{d}$ is a feature extractor that transforms a state into a $d$-dimensional vector, and $\boldsymbol{z}\in\mathbb{R}^{d}$ represents a latent vector. Randomized rewards for each state are generated by sampling $\boldsymbol{z}$ from a given distribution $P_{\boldsymbol{z}}$, and then setting rewards as $F(s,\boldsymbol{z})$. Even if $F(s,\boldsymbol{z})$ is high at unreachable parts of the state space, the agent simply does not collect the reward and moves on. This does not derail training as $\boldsymbol{z}$ is resampled at the start of each trajectory.
Latent conditioned policy and value network. Recall that the policy and the value function must be aware of the state and the random variable that factorizes the randomized reward function $F$. To achieve this, we augment the input to the policy and the value functions with the latent vector $\boldsymbol{z}$. The resulting policy and the value networks are $\pi(.|s,\boldsymbol{z})$ and $V^{\pi}(s,\boldsymbol{z})$. We train the latent-conditioned value network to approximate the expected sum of the original reward and the randomized rewards as below
$V^{\pi}(s,\boldsymbol{z})\approx\mathbb{E}_{\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t})+\lambda F(s_{t+1},\boldsymbol{z}))\right],$ (3)
where $\lambda$ is a hyperparameter. We train the latent-conditioned policy $\pi$ to maximize $V^{\pi}(s,\boldsymbol{z})$ at every state $s$ and latent vector $\boldsymbol{z}$. Both value and policy networks can be trained with any off-the-shelf RL algorithms, e.g. PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), A3C (Mnih et al., 2016), SAC (Haarnoja et al., 2018).
Latent vector sampling. To ensure the agent is exposed to a wide range of randomized reward functions, we randomize the latent vector $\boldsymbol{z}$, which is done by resampling at the start of each trajectory. This ensures each trajectory is rolled out under the same policy and latent vector $\boldsymbol{z}$, maintaining temporal consistency crucial for deep exploration, as indicated by prior work (Osband et al., 2016a; Fortunato et al., 2017). The sampling distribution of $\boldsymbol{z}$ is discussed in Section 4.
As we train the policy conditioned on the randomly sampled latent $\boldsymbol{z}$ that defines the randomized reward function, we term our method as Random Latent Exploration (RLE). We outline the algorithm in Algorithm 1 and present the detailed version in Algorithm 2 (in Appendix). Note that at line 6 in Algorithm 1, we compute the randomized reward using the next state $s_{t+1}$ since the next state reflects the effect of the agent’s chosen action $a_{t}$ in state $s_{t}$. This choice is common in prior works computing exploration bonuses, where the exploration bonus is a function of the newly-reached state rather than the current state (Burda et al., 2019).
@@ 4 Experiments
We compare RLE performance against the action noise exploration method typically used in many RL algorithms (Schulman et al., 2017; Mnih et al., 2015). In all our experiments, we train the agent using PPO (Schulman et al., 2017) for each task separately. Standard PPO implementation (Schulman et al., 2017) explores by sampling actions from the learned policy (i.e., a Boltzmann distribution over actions for discrete action spaces or a Gaussian distribution for continuous action spaces).
@@
  • NoisyNet (Fortunato et al., 2017): We chose it to be the representative baseline from the family of noise-based exploration (Osband et al., 2016a; Fortunato et al., 2017; Plappert et al., 2017) because it has been used in prior works on benchmarking exploration strategies and achieve superior performance(Chen et al., 2022; Taïga et al., 2019).
@@ 4.1 Illustrative Experiments on FourRoom
We first ran toy experiments on the FourRoom environment (Sutton et al., 1999) to test whether our method can perform deep exploration.
@@
Setup. Figure 1 illustrates FourRoom environment with $50\times 50$ states consisting of four rooms separated by solid walls (which the agent can’t cross) and connected with small openings of a single state each (which the agent needs to go through to travel across rooms). The agent observes the $(x,y)$ coordinates as the state and can take action to move left, right, up, and down (if not interrupted by a wall). At the beginning of each trajectory, the agent always starts from the top-right corner of the room (denoted by the letter “S”). In this study, we always give zero task reward to the agent since we are interested in comparing the deep exploration behavior of different strategies. This is also known as reward-free exploration (Pathak et al., 2017).We also perform experiments on FourRoom with an sparse task reward of $1$ at the bottom-left corner. The results and visualization of visitation counts are deferred to Appendix C.
Figure 1: FourRoom environment. The agent starts at the top-right state (denoted by red ‘S’) and can move left, right, up, and down. The black bars denote walls that block the agent’s movement.
@@
We compared agents trained with different exploration strategies: PPO, NoisyNet, RND, and RLE, over 2.5 million timesteps. For RLE, the feature extractor defined in Equation 2 is set to be a randomly initialized neural network with one hidden layer, the output layer of which has the same dimension as $\boldsymbol{z}\sim P_{\boldsymbol{z}}$ to make the final dot product operation feasible. Further implementation details are available in Section B.2. Because of the walls, the FourRoom environment requires deep exploration to go to states distant from the initial states.
@@
Does RLE qualitatively improve trajectory diversity during training? The exploration of RLE is driven by sampling the latent vector $\boldsymbol{z}$ to change the behavior of the policy network $\pi$ and the objective of the value network $V^{\pi}$. To investigate this, one may wonder how the choice of $\boldsymbol{z}$ affects the behaviors of the induced policy. To qualitatively understand this aspect, in Figure 2 we plotted different trajectories corresponding to different choices of the latent vector $\boldsymbol{z}\sim P_{z}$. Each trajectory corresponds to a specific choice of $\boldsymbol{z}$ and is assigned a unique color. For this plot, we chose the checkpoint of the policy network stored in the middle of training (i.e., 1.5 million timesteps) to observe the trajectory diversity of the agent as it explores during training. Figure 2 shows that the agent’s exploration is diverse and visits all four rooms. This simple experiment furthers our belief that diverse choices of latent vector $\boldsymbol{z}$ can induce diverse trajectories. Similar plots over more seeds for all the exploration algorithms RLE, PPO, RND, and NoisyNet are provided in Appendix C.
@@
Explaining the observed trajectory diversity. To explain why the trajectories generated by an RLE policy are diverse, we plot reward functions induced by sampling different $\boldsymbol{z}$ in Figure 10 (in Appendix). These plots demonstrate the diversity of the random rewards, each of which can guide the policy to a different part of the state space.
Quantifying State Visitation The state visitation counts of different exploration strategies is visualized in Figure 3. The results show that PPO’s state visitation centers around the starting room (i.e., top-right), indicating that action noise alone doesn’t encourage the agent to explore far from the initial state. In contrast, RLE, RND, and NoisyNet can frequently reach the rooms beyond the initial room, with RLE visitation count spread across the four rooms. This suggests that RLE can do deep exploration similar to prior deep exploration algorithms for this environment.
4.2 Benchmarking Results on Atari
Having performed illustrative experiments on the FourRoom toy environment, we now evaluate our method on more realistic and challenging tasks. Our results below show that RLE based exploration improves PPO’s overall performance on most tasks.
Setup. We evaluate our method in the well-known Atari benchmark (Bellemare et al., 2013). Following the common practice in Atari (Mnih et al., 2015), the agent observes a stack of the most recent four $84\times 84$ grayscale frames as input and outputs one of many discrete actions available in the Atari (see Bellemare et al. (2013) for further environment details). For RLE, we chose the feature learned by the value network followed by a randomly initialized linear layer as $\phi$ (used in Equation 2). Note that the randomly initialized linear layer is kept frozen throughout training. We set the dimension of the latent vector $\boldsymbol{z}$ as $8$. We test different values of the dimension of $\boldsymbol{z}$ in Section 4.4 on a subset of games and observe similar performance. We also test using a completely randomly initialized network as $\phi$ in Appendix D and observe slightly worse performance.
We use the standard PPO hyperparameters (Burda et al., 2019) for training, and all implementation details are provided in Appendix B. For each Atari game (i.e., environment), we train five policies with five different random seeds for $40$ million frames each following prior work (Chen et al., 2022; Bellemare et al., 2016). However, we trained Montezuma’s Revenge for $200$ million frames since its exploration difficulty is much harder than other Atari games (Burda et al., 2019).